OpenStudy (anonymous):

suppose y varies directly with x.if y=3 when x=2 find y when x=3

OpenStudy (anonymous):

y=kx find K

OpenStudy (anonymous):

i try that but still dont get it at all

OpenStudy (anonymous):

2=k*3 2/3 = k

OpenStudy (anonymous):

y = (2/3) 3

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

I did it opposite sorry. Its 3 = k*2 k = 3/2 y = (3/2) 3

OpenStudy (anonymous):

y = 4.5